# Personalized Generative Low-light Image Denoising and Enhancement

### v1

@@ 4.1 Dataset
Testing Data. For testing DiffPGD, we utilize both simulated and real-world low-light noisy data. The simulation setup assumes a $12$-bit image sensor with parameters similar to that of Canon DSLR cameras: a Full Well Capacity of $36000\text{e}^{-}$, Quantum Efficiency of $0.42$, dark current of $7\text{e}^{-}$, and read noise of $6\text{e}^{-}$. These parameters ensure realistic noise characteristics for evaluating model performance. We collect the real-world test data using the Sony $\alpha 6400$ DSLR camera at an approximate light intensity of $0.13$ Lux with an ISO of $25600$ and an exposure time of $\frac{1}{8}$ seconds.
@@ 4.2 Implementation Details
During the training stage, we train our model on the CelebAMask-HQ dataset [29], comprising 30,000 images distributed across 4,516 unique identities(IDs). We use 3318 IDs for training and 200 IDs for simulated testing. Within each ID, there are 3–20 images. The Adam optimizer is applied with a learning rate of 1e-4. Training is performed for 200,000 iterations with a batch size of 32, resulting in a total of 6,400,000 samples seen by the model.
Figure 6: Visual comparison on CelebAMask-HQ test set. Zoom in for the best visual experience.
Figure 7: Visual comparison on real-captured photo test (zoomed-in area in the input is scaled for visual purposes). Our method outputs the best denoised and enhanced result, and successfully restores fine facial details, such as moles, wrinkles.
@@ 4.3 Performance in Simulated Scenarios
In simulated scenarios, we compare our methods with several state-of-the-art works: UTVNet[65], ACENet[56], GFPGAN[50], SNR-aware[57], DR2[53], MIRNet[59] and DiffLL[21]. These works aim to restore high-quality images from degraded data that has been affected by low light, noise, or both, which are close to the subject of our research. Particularly, GFPGAN and DR2 are designed specifically for face image restoration, which is highly competitive with our method. We use the 200 test images which correspond to the 200 test IDs from CelebAMask-HQ and send them into our simulator to generate the test set. We simulated five low-light noisy levels by setting different photons per pixel (PPP), which range from 13 PPP to 65 PPP, the lower the PPP, severe the degradation is.
Quantitative and qualitative results are provided. Firstly, we use PSNR, SSIM, and the perceptual metric (LPIPS[62]) to measure the restored image quality. The results are shown in Table 1, it can be seen that our method has surpassed all previous methods by a large scale. Demonstrating the supreme ability of our proposed DiffPGD model on low light denoising and enhancement problems.
Since our work focuses on face images, we also check on the identity preservation performance. Therefore, besides the PSNR, SSIM, and LPIPS metrics, we adopt the identity score which calculates the cosine similarity of the image features extracted by ArcFace[10] to further evaluate the identity consistency. The results are shown in Table 2. Again, our method has the highest identity score. The advantage of our model on preserving the identity is especially outstanding in heavily degraded cases.
Visualized comparisons are presented in Fig. 6. The image restored from our method has the highest quality and more details, meanwhile preserving the facial identity and maintaining the ID consistent best. Again, the visual results are consistent with the quantitative metric results, the advantage of our model on preserving the identity is especially outstanding in the heavily degraded cases. The sampled gallery photos of the test subject can be found in Fig. 19. Refer to Sec. B in the supplementary material for more visual results.
  Photons Per Pixel 13 26 39 52 65
[…]
UTVNet[65] 12.08 0.2644 0.7816 13.26 0.3685 0.7169 13.99 0.4332 0.6754 14.50 0.4783 0.6419 14.87 0.5118 0.6148
ACENet[56] 11.25 0.2815 0.7431 12.27 0.4022 0.6725 13.19 0.4670 0.6276 13.87 0.5120 0.5894 14.43 0.5479 0.5577
SNR-aware[57] 10.74 0.1186 0.8193 11.99 0.2031 0.7650 12.34 0.2546 0.7230 12.53 0.2935 0.6852 12.76 0.3271 0.6507
MIRNet[59] 11.64 0.4553 0.7476 13.64 0.5416 0.6491 13.90 0.5724 0.5874 14.12 0.5960 0.5355 14.323 0.6154 0.4934
DiffLL[21] 13.30 0.3472 0.7872 14.66 0.4282 0.7189 15.47 0.4588 0.6736 15.96 0.4805 0.6387 16.25 0.4969 0.6072
DR2[53] 13.50 0.3448 0.7573 15.16 0.5141 0.6430 16.22 0.5655 0.5871 16.85 0.5909 0.5503 17.54 0.6209 0.5181
GFPGAN[50] 13.53 0.4256 0.7049 15.56 0.5446 0.5713 16.60 0.5960 0.5156 17.41 0.6322 0.4748 17.79 0.6563 0.4431
Ours 21.10 0.6287 0.3862 22.80 0.6914 0.3394 23.85 0.7195 0.3156 24.11 0.7354 0.2965 24.38 0.7399 0.2947
Table 1: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset.
  Methods 13 PPP 26 PPP 39 PPP 52 PPP 65 PPP
UTVNet[65] 0.1572 0.3654 0.5201 0.6197 0.6938
ACENet[56] 0.1734 0.4173 0.5769 0.6756 0.7240
SNR-aware[57] 0.1069 0.1921 0.3114 0.4171 0.4971
MIRNet[59] 0.3071 0.5613 0.6694 0.7246 0.7794
DiffLL[21] 0.1751 0.3779 0.5033 0.6107 0.6682
DR2[53] 0.2456 0.5128 0.6358 0.6905 0.7239
GFPGAN[50] 0.2506 0.5392 0.6569 0.7212 0.7615
Ours 0.5995 0.7271 0.7702 0.8044 0.8253
Table 2: Comparisons of identity score with other methods on CelebAMask-HQ dataset.
UTVNet[65] 12.02 0.3901 0.6954 0.8062
ACENet[56] 11.36 0.3962 0.6817 0.8125
SNR-aware[57] 10.18 0.1922 0.7274 0.6111
MIRNet[59] 12.48 0.4315 0.7132 0.6914
DiffLL[21] 13.18 0.3899 0.7551 0.7686
DR2[53] 14.08 0.1829 0.7350 0.5309
GFPGAN[50] 14.18 0.3714 0.6784 0.7138
Ours 14.02 0.4362 0.5954 0.8248
@@
Table 3: Comparisons with state-of-the-art works on real-captured photo test.
w/o physical buffers 12.96 0.3801 0.6196 0.7538
w/ physical buffers 14.02 0.4362 0.5954 0.8248
@@
4.4 Performance in Real-Life Scenarios
Testing model performance in real-life low-light scenarios is crucial, as real-world conditions often involve uncontrollable factors compared to simulated ones. Since no existing dataset addresses the restoration of real-life low-light noisy face photos, we created a custom test set. Specifically, we printed five photos of five target individuals and captured images of these photos under low-light conditions using a camera with the settings specified in Section 4.1. As shown in Fig. 7, our model demonstrates significantly better recovery of real-captured low-light noisy inputs compared to existing models. In Table 3, we quantitatively compare the performance of our model with other baselines on this real-captured photo test. We observe that all our metrics are better except for GFPGAN’s PSNR. However, when looking at the visual comparison in Fig. 7, we realize that GFPGAN recovers smoother features but loses the true facial details, resulting in a relatively fake-looking face. For further verification, we also photographed real individuals under low-light conditions with the camera. As shown in Fig. 8, DiffPGD outperforms existing methods by a large margin. These results confirm that DiffPGD effectively leverages prior information from a target person’s gallery photos to guide the diffusion model in learning a personalized generative space. During restoration, the model accurately directs the output toward this clean personalized space. The gallery photos of the real-captured male subject are displayed in Fig. 9. Other test subjects’ gallery photos can be found in Fig. 19.
Figure 8: (Left) Environmental setup, real-life low-light noisy photo taken in the dark room. (Right) Visual comparison on the real-world captured SUBJECT.
Figure 9: Male SUBJECT’s gallery (face areas are cropped from the subject’s photos).
@@ 5 Ablation Study
ID-Consistent Physical Buffers for Identity Preservation. We have shown that the proposed model DiffPGD exhibits outstanding performance in both simulated and real-world testing scenarios. In Table 4, we compare the quantitative results of training our model with and without ID physical buffers in real-world test cases. The results demonstrate that using physical buffers largely increases the identity score, highlighting the effectiveness of extracting ID physical buffers from a person’s existing photo gallery to assist in restoring real-world low-quality photos.
@@ 6 Conclusion
In this work, we addressed the challenges of low-light image restoration by introducing Personalized Generative Denoising (PGD), a novel approach that leverages user-specific photo galleries to improve denoising performance. By incorporating an identity-consistent physical buffer, we provided a strong prior that enhances the diffusion model’s ability to restore degraded images while avoiding the need for model fine-tuning. Our experiments demonstrated that PGD outperforms existing diffusion-based denoising methods, delivering superior results across various testing scenarios. This approach not only mitigates false content generation in low-SNR conditions but also bridges the gap in low-light imaging quality, paving the way for personalized and more robust image restoration techniques.
Acknowledgment This research was supported, in part, by a research contract from Samsung Research America, the National Science Foundation (NSF), and in part by the Intelligence Advanced Research Projects Activity (IARPA) under Contract No. 2022-21102100004, and in part by the National Science Foundation under the grants CCSS-2030570 and IIS-2133032. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix B Additional Visual Results
In Fig. 16, Fig. 17 and Fig.18, we present additional visual comparisons between our model and baseline methods. Our results consistently demonstrate superior performance, particularly in cases of severe degradation.
Appendix C Results in More Severe Cases
To further evaluate the ability of our model to denoise and enhance extremely degraded photos, we tested it under highly challenging conditions, including both simulated and real-life scenarios.
For simulated cases, We simulate severely degraded test cases with 10 PPP and 5 PPP. As shown in Fig.12 and Fig.13, our model can still perform well in terms of enhancing and denoising the degraded inputs, whereas other models struggled significantly. Besides, due to the extremely low signal level in the input image, we can see some hallucination effect in the hairstyle of the generated outputs. Despite this, thanks to the ID physical buffer extracted from the gallery of the target person, our model could still effectively restore the facial structure of the target person. Quantitative comparisons are presented in Tables 7 and 8.
We also evaluated the model on real-captured photos. Fig. 14 is a visual example of the male subject taken under extremely low light conditions with high noise levels. From the figure, we can see that our proposed model outperforms all baseline methods in both denoising and enhancement. Additionally, the blind-based restoration methods DR2 and GFPGAN completely failed to produce meaningful results due to the severely low signal levels. This supports our observation that blind restoration approaches struggle with images degraded by extremely low light and noise. For these real-captured extreme test cases, we further conducted quantitative evaluations of both image quality and identity preservation, as summarized in Table 6. Since ground truth images are unavailable for real-captured cases, we computed the Fréchet Inception Distance(FID)[19] between the model’s output images and the subject’s gallery photos, which are 4 clean natural photos from the subject’s phone (shown in Fig.9). We calculate the identity (ID) score by averaging the ID scores between the model’s output images and each gallery photo. The results indicate that our model consistently achieves superior image quality and identity preservation compared to other methods.
Figure 10: The relationship between the identity (ID) score of the restored images and the number of gallery images used to generate the ID physical buffer.
@@ Appendix C Results in More Severe Cases
  Methods FID $\downarrow$ ID Score $\uparrow$
UTVNet[65] 293.2 0.31
ACENet[56] 338.5 0.13
MIRNet[59] 387.7 0.13
DiffLL[21] 355.1 0.36
DR2[53] 341.0 0.06
GFPGAN[50] 316.8 0.06
Ours 284.2 0.43
Table 6: Comparisons with state-of-the-art works on the real-captured test cases under extreme conditions.
  Photons Per Pixel 5 10
[…]
UTVNet[65] 11.06 0.1756 0.8275 11.71 0.2226 0.8067
ACENet[56] 9.97 0.1741 0.7949 10.81 0.2326 0.7683
SNR-aware[57] 9.11 0.0590 0.8527 9.91 0.0864 0.8366
MIRNet[59] 9.05 0.3376 0.8546 10.02 0.3896 0.8004
DiffLL[21] 11.96 0.2832 0.8448 12.67 0.3109 0.8143
DR2[53] 7.17 0.1902 0.7435 8.08 0.2399 0.7437
GFPGAN[50] 7.20 0.2211 0.7251 8.13 0.2845 0.7185
Ours 17.93 0.5375 0.4731 18.49 0.6115 0.4120
Table 7: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset under extreme conditions.
  Methods 5 PPP 10 PPP
UTVNet[65] 0.0998 0.1129
ACENet[56] 0.0999 0.1233
SNR-aware[57] 0.0988 0.1001
MIRNet[59] 0.0991 0.1747
DiffLL[21] 0.1004 0.1209
DR2[53] 0.0987 0.1043
GFPGAN[50] 0.0949 0.1176
Ours 0.5995 0.7271
Table 8: Comparisons of identity score with other methods on CelebAMask-HQ dataset under extreme conditions.
@@
Table 9: Comparison of ID physical buffers and encoder-based methods on the real-world photo test set. For a fair comparison, we use two same gallery images for each individual for both the encoder-based and physical buffer-based approaches.
Appendix D More Ablation Study
The proposed DiffPGD model is conditioned with the ID physical buffers extracted from the User’s gallery photos. As claimed in the paper, facial features extracted from a single image are inherently limited. In Fig. 10, we evaluate the effect of the number of gallery images used to generate the ID physical buffer on identity preservation performance. We analyze it with the real-captured extreme test set where 4 gallery images were collected from the subject’s phone. The results show that identity preservation improves as the ID physical buffer is derived from a greater number of gallery photos rather than a single image. Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it, which corresponds to the case with zero gallery photos.
Appendix E More Implementation Details
During training and testing, the image size is 256 $\times$ 256. The architecture of the diffusion model is adapted from [11]. We modify the architecture so that the model can take the concatenation of physical buffers and degraded images as conditions. The model parameters are configured as follows: the number of channels is set to 128, channel multiple is (1, 1, 2, 2, 4, 4), head channels are 128, attention resolution is 16, dropout rate is 0.1, diffusion steps are 1000, and the two hyperparameters P2_gamma and p2_k are both set to 1.0. For additional architectural details, refer to [11]. During inference, we adopt the DDIM method [45] with 200 denoising steps. The model is trained using the Adam optimizer [26] with a learning rate of $10^{-4}$. The weight decay is set to 0. All experiments are performed on a single NVIDIA A100 GPU.
Appendix F Encoder vs. Physical Buffer?

### v2

@@ 4 Experiments
Figure 6: Visual comparison on real cases. The first column presents the input images along with their capture parameters. In the second column, we enhance the brightness to better reveal the noise. Please zoom in for the best visual experience.
@@ 4.1 Dataset
  Photons Per Pixel 5 10 13 26
Methods ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$
SNR-aware[59] 0.2156 188.3311 0.1302 8.4541 0.4240 139.8193 0.0786 9.5462 0.4999 124.1162 0.0626 10.1712 0.6937 99.9875 0.0383 12.4961
MIRNet[61, 62] 0.1229 294.7381 0.2414 8.6008 0.2860 214.7785 0.1534 9.8572 0.3830 187.2571 0.1286 10.4209 0.6212 136.1320 0.0793 12.8992
DiffLL[22] 0.2910 167.3905 0.1068 17.3506 0.4544 150.4559 0.1117 18.3360 0.5095 140.2738 0.0996 19.9547 0.6833 119.2693 0.0760 23.3180
GFPGAN[52] 0.1941 388.9293 0.4256 7.1990 0.2075 233.3173 0.1521 8.1283 0.2976 190.7264 0.1326 9.5483 0.5976 158.9475 0.0829 11.4855
CodeFormer[67] 0.2863 111.4544 0.04908 18.2406 0.4940 91.1244 0.0326 19.2692 0.5563 84.7955 0.0281 20.2306 0.7338 80.6582 0.0263 24.5512
Ours 0.3714 77.5743 0.0084 17.9334 0.5225 71.8645 0.0063 18.4999 0.5938 69.5567 0.0042 19.0910 0.7118 59.1975 0.0030 23.7226
Table 1: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset on 5PPP, 10PPP, 13PPP, and 26PPP. The lower the PPP, severe the degradation is.
Testing Data. For testing DiffPGD, we utilize both simulated and real-world captured low-light noisy data.
The simulation setup assumes a $12$-bit image sensor with parameters similar to that of Canon DSLR cameras: a Full Well Capacity of $36000\text{e}^{-}$, Quantum Efficiency of $0.42$, dark current of $7\text{e}^{-}$, and read noise of $6\text{e}^{-}$. These parameters ensure realistic noise characteristics for evaluating model performance.
We also captured 100 RAW-format face images as real-world test dataset using a Sony $\alpha 6400$ DSLR camera paired with a Sony 18-135mm lens at night. These images feature diverse low-light and dark backgrounds, various angles and poses, as well as different focal lengths, apertures, ISO values, and shutter speeds to reflect the rich characteristics of real-world low-light photography. We cropped these images into 256×256 face patches.
@@ 4.2 Implementation Details
During the training stage, we train our model on the CelebAMask-HQ dataset [30], comprising 30,000 images distributed across 4,516 unique identities(IDs). We use 3318 IDs for training and 200 IDs for simulated testing. Within each ID, there are 3–20 images. The number of gallery images we used per person varies from 3 to 6, with most around 5. The Adam optimizer is applied with a learning rate of 1e-4. Training is performed for 200,000 iterations with a batch size of 32, resulting in a total of 6,400,000 samples seen by the model.
4.3 Performance in Real-Captured Scenarios
Testing model performance in real-life low-light scenarios is crucial and challenging, as real-world conditions often involve uncontrollable factors compared to simulated ones. Since no existing data set addresses the restoration of real-life low-light noisy face photos, we captured photos of real persons under indoor and outdoor low-light scenarios , as illustrated in Section 4.1.
We compare our methods with several state-of-the-art works: UTV-Net[66], GFPGAN[52], SNR-aware[59], MIRNet[61, 62], DiffLL[22], Codeformer [67]. These works aim to restore high-quality images from degraded data that has been affected by low light, noise, or both, which are close to the subject of our research. Particularly, GFPGAN and Codeformer are designed specifically for face image restoration, which is highly competitive with our method. We re-trained/fine-tuned all the baselines with the same training dataset as ours for fair comparison.
Quantitative and qualitative results are presented. To assess the quality of restored images, we employ the Fréchet Inception Distance (FID) [20] and the Kernel Inception Distance (KID) [1], as these metrics offer a robust evaluation aligned with human perception. Since ground-truth images are unavailable for real-captured cases, we compute the FID and KID score between the model’s output and the corresponding subjects’ gallery photos. As shown in Fig. 6, our method produces images with the highest quality and level of detail. The restored images exhibit more natural skin tones and well-preserved facial identity features, such as eye shape, eyebrows, and lips, while maintaining the highest consistency in facial identity.
Since our work focuses on face images, we also check on the identity preservation performance. Besides the image quality metrics, we adopt the identity score which calculates the cosine similarity of the image features extracted by ArcFace[11] to further evaluate the identity consistency. Again, since ground truth images are unavailable for real-captured cases, we calculate the identity (ID) score by averaging the ID scores between the model’s output images and each gallery photo of the corresponding subjects. The results are also shown in Table 2. Again, our method has the highest identity score, and the visual results are consistent with the quantitative metric results. The results demonstrate the supreme ability of our proposed DiffPGD model on real low light denoising and enhancement task.
@@ 4.4 Performance in Simulated Scenarios
In simulated scenarios, We use the 200 test images which correspond to the 200 test IDs from CelebAMask-HQ and send them into our simulator to generate the test set. We simulated various low-light noisy levels by setting different photons per pixel (PPP), i.e., the average number of photons arriving at the sensor within the exposure time. The lower the PPP, severe the degradation is. A lower photon level increases the impact of read noise and dark current on the final output, thereby decreasing the quality of the reconstructed images.
Quantitative results are provided in Table 1. The image restored from our method has the highest quality and more details, meanwhile preserving the facial identity and maintaining the ID consistent best. (Refer to supplementary for more visual comparisons). From the table, we could see that CodeFormer surpasses us on PSNR. Fig. 7 illustrates a visual comparison between our method and CodeFormer under inputs ranging from 5 to 65 PPP. We observe that CodeFormer tends to generate smoother images while we could preserve better high-frenquency details. Furthermore, our method demonstrates superior identity preservation even at low resolutions, such as 5 PPP to 13 PPP, while exhibiting fewer artifacts. The advantage of our model in maintaining identity consistency is particularly evident in heavily degraded cases. The results highlight the strong performance of our proposed DiffPGD model in low-light denoising and enhancement tasks.
  Methods FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$
SNR-aware[55] 252.2397 0.1264 0.1990
UTVNet[66] 341.8502 0.1617 0.2002
MIRNet[61, 62] 276.3853 0.1659 0.2210
DiffLL[22] 252.1553 0.1421 0.3196
GFPGAN[52] 252.7411 0.1249 0.1525
CodeFormer[67] 281.9744 0.1459 0.2965
Ours 251.9318 0.0882 0.3708
@@
Table 2: Comparisons with state-of-the-art works on the real-captured test cases.
  Methods FID $\downarrow$ KID $\downarrow$ ID Score $\uparrow$
w/o physical buffers 272.6148 0.1281 0.3169
w/ physical buffers 251.9318 0.0882 0.3708
@@
Figure 7: Visual comparison on the simulated cases. Our method maintains good identity consistency even at extremely low PPP and exhibits fewer artifacts.
Figure 8: The relationship between the identity (ID) score and FID of the restored images and the number of gallery images used to generate the ID physical buffer.
@@ 5 Ablation Study
ID-Consistent Physical Buffers for Identity Preservation. We have shown that the proposed model DiffPGD exhibits outstanding performance in both simulated and real-captured testing scenarios. In Table 3, we compare the quantitative results of training our model with and without ID physical buffers in real-captured test cases. The results demonstrate that using physical buffers largely increases the identity score, highlighting the effectiveness of extracting ID physical buffers from a person’s existing photo gallery to assist in restoring real-world low-quality photos.
Effect of the Number of Gallery Images. The proposed DiffPGD model is conditioned with the ID physical buffers extracted from the User’s gallery photos. As claimed in the paper, facial features extracted from a single image are inherently limited. In Fig. 8, we evaluate the effect of the number of gallery images used to generate the ID physical buffer. We analyze it with the real-captured set where 6 gallery images were collected from each subject. The results show that identity preservation improves as the ID physical buffer is derived from a greater number of gallery photos rather than a single image. Additionally, incorporating the ID physical buffer significantly enhances identity preservation compared to not using it, which corresponds to the case without gallery photos.
@@ 6 Conclusion
In this work, we addressed the challenges of low-light facial image restoration by introducing Diffusion-based Personalized Generative Denoising (DiffPGD), a novel approach that leverages user-specific photo galleries to improve denoising performance. By incorporating an identity-consistent physical buffer, we provided a strong prior that enhances the diffusion model’s ability to restore degraded images while avoiding the need for model fine-tuning. Our experiments, which included extensive evaluations on multiple challenging datasets, demonstrated that DiffPGD not only outperforms existing diffusion-based denoising methods but also delivers superior quantitative and qualitative results across various testing scenarios. This approach effectively mitigates false content generation in low-SNR conditions while bridging the gap in low-light imaging quality, ultimately paving the way for personalized and more robust image restoration techniques.
@@ Appendix B Additional Visual Results
In Fig. 12, Fig. 14, and Fig.15, we present additional visual comparisons between our model and baseline methods on real-captured and simulated cases. Our results consistently demonstrate superior performance, particularly in cases of severe degradation.
@@
Table 6: Comparison of ID physical buffers and encoder-based methods on the real-captured photo test set. For a fair comparison, we use two same gallery images for each individual for both the encoder-based and physical buffer-based approaches.
Appendix C More Implementation Details
During training and testing, the image size is 256 $\times$ 256. The architecture of the diffusion model is adapted from [12]. We modify the architecture so that the model can take the concatenation of physical buffers and degraded images as conditions. The model parameters are configured as follows: the number of channels is set to 128, channel multiple is (1, 1, 2, 2, 4, 4), head channels are 128, attention resolution is 16, dropout rate is 0.1, diffusion steps are 1000, and the two hyperparameters P2_gamma and p2_k are both set to 1.0. For additional architectural details, refer to [12]. During inference, we adopt the DDIM method [46] with 100 denoising steps. The model is trained using the Adam optimizer [27] with a learning rate of $10^{-4}$. The weight decay is set to 0. All experiments are performed on a single NVIDIA A100 GPU.
Appendix D Encoder vs. Physical Buffer?